An English class consists of twenty students, and five are to be chosen to give speeches in a school competition. In how many different ways can the teacher choose the team, given the following conditions? (a) The order of the speakers is important. 7296 Incorrect: Your answer is incorrect. (b) The order of the speakers is not important.
Answer #1

Answer)

When order is not important we use combinations

Ncr = n!/(r!*(n-r)!)

Where n! = n*n-1*n-2*n-3*n-4*n-5*....till 1

For example

5! = 5*4! = 5*4*3! = 5*4*3*2! = 5*4*3*2*1

When order is important we use permutations

Npr = n!(n-r)!

A)

Order is important

N = 20

R = 5

20p5 = 20!/(20-5)! = 20!/15! = 1860480

B)

Order is not important

20c5 = 20!/(5!*15!) = 15504